# Tuning the Frequencies: Robust Training for Sinusoidal Neural Networks

### v2

@@ 1 Introduction
Sinusoidal multilayer perceptrons (MLPs) emerged as powerful implicit neural representations (INRs) for low-dim signals [27, 3, 35, 26, 7, 22, 9, 10, 8, 34]. In this context, the INR $f$ should fit the input data $\{x_{i},\mathscr{f}_{i}\}$ as close as possible, i.e. $f(x_{i})\approx\mathscr{f}_{i}$, without overfitting, thus encoding the signal implicitly in the MLP parameters. Therefore, two major properties are required: (1) $f$ needs high representation capacity to fit $\{x_{i},\mathscr{f}_{i}\}$; (2) $f$ should have bandlimit control to avoid frequencies bypassing the sampling rate.
@@
Overall, initializing and training sinusoidal MLPs remain empirical tasks which lack on deeper understanding to guide the learning process. To approach these problems, we present TUNER (TUNing sinusoidal nEtwoRks), a training scheme for sinusoidal MLPs consisting of a robust initialization and bandlimit mechanism that greatly improves the capacity and convergence of INRs. In contrast to previous work, TUNER is grounded on a theoretical framework that guides our learning approach. We develop a new trigonometric identity (Thrm 1), resembling a Fourier series, and prove that such expansion produces a large number of frequencies in terms of the first layer (input frequencies). This explains the frequency generation process when composing layers, highlighting the importance of input layer initialization for improving capacity. Remarkably, the amplitudes of these frequencies are given by complex functions of the hidden weights, introducing a challenging training. To address this, we prove that those are bounded by a term depending only on the hidden weights (Thrm 2). We apply this result to control the bandlimit of the INR during training. Some of TUNER’s results are showcased in Fig. 1, in comparison to previous work, illustrating substantial enhancements to sinusoidal INR modeling. In summary, our contributions are:
  • We introduce a novel trigonometric identity that expands each sinusoidal neuron into a wide sum of sines with frequencies being integer combinations of the input frequencies (Fig 2). The corresponding amplitudes, determined by the hidden weights, have a useful upper bound, offering effective tools for controlling the resulting signal.
@@ 2 Related works
Implicit neural representations (INRs) [31] are a trending topic in machine learning, used to learn highly detailed signals in low-dimension domains. Current INR architectures use Fourier feature mappings [29] or sinusoidal activation functions [27] to bypass the spectral bias [19] common in ReLU multilayer perceptrons. The high representation capacity of sinusoidal INRs has motivated their use to represent a wide range of media objects. Examples include audio [6, 27], images [4, 28], face morphing [23], signed distance functions [24, 13, 27, 32, 15], displacement fields [33], surface animation [14, 16], multi-resolution signals [21, 13, 5, 30, 18], among others. Most of these exploit the derivatives of sinusoidal INRs in the loss functions.
@@ 3 Sinusoidal MLPs as Fourier series
Sinusoidal MLPs demonstrated high representational capacity with only a few hidden layers [27, 35, 21]. To understand how layer composition encodes this capacity, we propose to thoroughly investigate the structure of a sinusoidal MLP $f\!:\!\mathbb{R}^{d}\!\to\!\mathbb{R}$. For simplicity, we assume the codomain to be $\mathbb{R}$, however the same analysis holds for dimension $>\!\!1$. We assume $f$ has a single hidden layer $\textbf{S}\!:\!\mathbb{R}^{m}\!\to\!\mathbb{R}^{n}$, with $m,n\!\in\!\mathbb{N}$. Precisely, we consider $f(\textbf{x})\!=\!\textbf{C}\!\circ\!\textbf{S}\!\circ\!\textbf{D}(\textbf{x})\!+\!e,$ with $\textbf{D}(\textbf{x})\!=\!\sin(\omega\textbf{x}\!+\!\varphi)$ being the input layer that projects $\textbf{x}\in\!\mathbb{R}^{d}$ into a list of harmonics (input neurons) ${D}_{i}(\textbf{x})\!=\!\sin\left(\omega_{i}\textbf{x}+\varphi_{i}\right)$ with frequencies $\omega\!=\!(\omega_{1},\ldots,\omega_{m})\!\in\!\mathbb{R}^{m\times d}$ and shifts $\varphi\!\in\!\mathbb{R}^{m}$. Layer D is then composed with $\textbf{S}(\textbf{x})\!=\!\sin(\textbf{W}\textbf{x}+\textbf{b})$, where $\textbf{W}\in\mathbb{R}^{n\times m}$ is the hidden matrix, and $\textbf{b}\!\in\!\mathbb{R}^{n}$ the bias. Finally, $\textbf{C}\cdot\textbf{x}+e$ is an affine transformation with $\textbf{C}\in\mathbb{R}^{n\times 1}$ and $e\in\mathbb{R}$.
@@
Each neuron $h_{i}$ of a sinusoidal MLP has an amplitude-phase expansion of the form
@@
The Bessel functions $J_{k}$ appear in the Fourier series of $\sin\big{(}a\sin(x)\big{)}$ [2, Page 361] and Thrm 1 generalizes this result. The proof, details, and a generalization to deeper neurons can be found in the Supp. Mat.
@@ 4.1 Initialization as spectral sampling
Figure 4: Uniform initialization of $\omega$ (top), as in SIREN, and our initialization (bottom). Ours offers better signal/gradient reconstruction w/o gradient supervision. The MLPs $(m\!=\!128,n\!=\!256)$ with bandlimit $\mathscr{b}\!=\!87$ were trained during $3000$ epochs.
@@ 4.2 Bounding and hidden layer initialization
We use these bounds to initialize W such that each column is normal distributed with mean 00 and standard deviation $\tfrac{\mathscr{c}_{\textbf{L}}}{3}$, with $\mathscr{c}_{\textbf{L}}$ being the bound corresponding to that column. Note that SIREN [27] initializes $\textbf{W}\in\mathbb{R}^{n\times m}$ uniformly in $\small{\big{(}\!\!-\!{\sqrt{{6}/{m}}},{\sqrt{{6}/{m}}}\big{)}}$, and as $\small{\left|\textbf{W}\right|_{\infty}\!\leq\!\sqrt{{6}/{m}}<2}$ for $m>1$, it is bandlimited only at the beginning of training, since W is not controlled in any way during the optimization. Conversely, our initialization controls the bandlimit during all training and has faster convergence.
Importantly, we present a scheme to learn the bounding parameters during training. For this, we bound each $j$-column of the hidden matrix W using a $\tanh$ activation followed by multiplication with a trainable bounding parameter $\mathscr{c}_{j}$. Clearly, the resulting column is bounded by $\mathscr{c}_{j}$. In other words, we replace the hidden layer $\sin(\textbf{W}\textbf{x}+\textbf{b})$ by
@@ 5 Experiments
This section presents experiments regarding the initialization and bounding of frequencies of an INR $f$ considering the images from Kodak dataset [1] resized to $[-1,1]^{2}$. Again, we describe the architecture of $f$ with hidden matrix $\textbf{W}\in\mathbb{R}^{n\times m}$ with the parameters $m,n$. We initialize the network using the scheme given in Sec 4.1. We use Adam’s optimizer [12] with learning rate $41e\!-\!$ 4. For the quantitative evaluations, we report the mean and standard deviation of the Peak Signal-to-Noise Ratio (PSNR) evaluated on the test set ($10\%$ random pixels). We compare the reconstruction quality of our INR $f$ and the resulting analytical gradient $\nabla f$ to show that our method adds high order regularity to the reconstruction. All experiments were performed in a GPU NVIDIA TITAN X (Pascal) with 12 GB.
@@ 5.1 Initialization
First, we compare our initialization of $\omega$ against the uniform distribution. We obtain a 4dB improvement of PNSR on average (with a std of 1.6dB). We also present an ablation over the parameter $\mathscr{l}$ which splits the square $[-\mathscr{b},\mathscr{b}]^{2}$ in regions of low (L) and high (H) frequencies. Here we use $\mathscr{b}=85$ and $\mathscr{l}\!=\!21$, $42$, $64$, and sample $\omega$ with $30\%$, $50\%$, and $70\%$ of coordinates in H. Table 1 shows the performance of the INR, with the signal (grad.) mean std of $1.9$dB ($3.1$dB).
$30\%$ $50\%$ $70\%$
@@
Figure 8: Reconstructions with large $\mathscr{l}$ present artifacts (overfitting) on the $10\%$ unsupervised points.
@@ 5.2 Bounding
Fig 9 compares the reconstructions of an image of size $512^{2}$ wo/w our bounding scheme. To demonstrate the impact of our scheme, we sample the input frequencies using $\mathscr{b}\!=\!256$ and $\mathscr{l}\!=\!100$. Without bounding (left), we effectively reconstruct the signal at the supervised pixels (30.8 dB), but with a noisy gradient (15.8 dB). That is, we are overfitting the signal. Conversely, bounding the hidden weights significantly improves both the signal (35.1 dB) and its gradient (27.8 dB). Our bounding scheme acts like a filter in both signal and gradient reconstruction.
Figure 9: Comparison of training wo/w bounds. Note that training using bounds preserves high order information of the signal.
We compare quantitatively our bounding scheme over the dataset, varying the parameters $\mathscr{b}$ and $\mathscr{c}_{\textbf{H}}$. Table 2 shows an improvement in the signal for all cases. Furthermore, the unsupervised gradient of the image is better learned when using bounds, showing an average increase of 9.6dB against the standard training. Also, the last two lines of the table show the impact of the parameters $\mathscr{c}_{\textbf{L}}$, $\mathscr{c}_{\textbf{H}}$ over the signal and gradient reconstructions. We consider training with $\mathscr{c}_{\textbf{L}}>\mathscr{c}_{\textbf{H}}$ since we are prioritizing low input frequencies in the spectrum generation. Otherwise, this may generate high frequencies, possibly bypassing the Nyquist limit and resulting in a worse gradient reconstruction.
@@
Table 2: Training with bounding improves the signal and grad. reconstruction. The choice of $\mathscr{c}_{\textbf{H}}$ greatly impacts the grad. fidelity. The std of the difference between each image signal (grad) quality wo/w bounding is $1.4$dB (1.6dB) on average. Evaluations performed on the test set ($10\%$ of pixels).
@@
Figure 10: Comparison of training with fixed/learned bounds. Learning the bound of each frequency allows an adjustment of the signal spectrum resulting in better signal/gradient reconstructions. Each blue point in the diagram represents a pair $(\omega_{j},\mathscr{c}_{j})$.
As expected, training lowers the bounds for high frequencies (H) while maintaining or increasing the bounds of low frequencies (L). Thus, we achieve better reconstruction in both the signal and gradient. In Fig 10 (right), we note that low frequencies (left of the green line) have significant bounds, showing their importance during training. These schemes allow us to initialize with higher frequencies than $\nicefrac{{B}}{{3}}$, accelerating convergence.
@@ 5.3 Comparisons
We first compare our method (TUNER) with SIREN, which requires choosing a frequency parameter $\mathscr{b}$, [^2] Denoted as $\omega_{0}$ in [27]. a critical task since higher $\mathscr{b}$ implies noisy reconstructions. TUNER also uses $\mathscr{b}$ to control the frequency generation, however, it provides more stable training with better reconstruction. To demonstrate this, we perform a comparison by varying $\mathscr{b}$ and the number of input ($m$) and hidden ($n$) neurons. First, consider the case where we have no information about the signal bandlimit besides the Nyquist limit $B$. Fig 11 presents a comparison between SIREN (top) and TUNER (bottom) with $\mathscr{b}=B$. Note that SIREN produces noisy reconstructions, while ours learns faster and does not suffer from noise. This is due to our bounding scheme, which limits the generation of high frequencies and provides stable training (Sec 4.2). In contrast, SIREN initializes its parameters such that the spectrum is initially bounded but does not maintain this guarantee during training.
Figure 11: Comparing SIREN (top) with TUNER (bottom) when we only know the Nyquist limit. Note that SIREN introduces high frequencies at the beginning of training, resulting in very noisy reconstructions. Conversely, ours starts training low frequencies, obtaining a faster convergence.
@@
Figure 12: Comparison with SIREN considering $\mathscr{b}=40$. Our method improves training convergence in both cases ($m<n$ and $m>n$) and better training stability when $m>n$.
Next, we numerically evaluate these experiments, providing a quantitative comparison with SIREN (and BACON) on the Kodak dataset. We compare different architectures ($m$, $n$) and spectrum bandlimits $\mathscr{b}$, evaluating the results in terms of signal and gradient PSNR over the supervised (Train) and unsupervised (Test) pixels. Table 3 presents these evaluations.
@@
As expected, SIREN performs well with $\mathscr{b}=\frac{B}{3}$. In this case, our method is perceptually similar to SIREN but improves gradient reconstruction and consistently outperforms BACON. When $\mathscr{b}=171$, our method demonstrates the best quality in both signal and gradient, showing greater robustness to spectral noise and superior performance even with $\mathscr{b}\geq\nicefrac{{B}}{{3}}$. Finally, with $\mathscr{b}=256$, our method surpasses SIREN in all cases. For BACON, our method achieves better signal reconstruction, with comparable gradients. However, the following experiment shows that BACON’s reconstruction has visible defects in this scenario. Note that we do not supervise the training with gradient information.

### v3

@@ 1 Introduction
Sinusoidal multilayer perceptrons (MLPs) emerged as powerful implicit neural representations (INRs) for low-dimensional signals [26, 3, 35, 11]. In this context, the INR $f$ should fit the input data $\{x_{i},\mathscr{f}_{i}\}$ as close as possible, i.e. $f(x_{i})\approx\mathscr{f}_{i}$, without overfitting, thus encoding the signal implicitly in the MLP parameters. Therefore, two major properties are required: (1) $f$ needs high representation capacity to fit $\{x_{i},\mathscr{f}_{i}\}$; (2) $f$ should have bandlimit control to avoid frequencies bypassing the sampling rate.
@@
Overall, initializing and training sinusoidal MLPs remain empirical tasks which lack on deeper understanding to guide the learning. To approach these problems, we present TUNER (TUNing sinusoidal nEtwoRks), a training scheme for sinusoidal MLPs consisting of a robust initialization and bandlimit mechanism that greatly improves the capacity and convergence of INRs. In contrast to previous work, TUNER is grounded on a theoretical framework that guides our learning approach. We develop a new trigonometric identity (Thrm 1), resembling a Fourier series, and prove that such expansion produces a large number of frequencies in terms of the first layer (input frequencies). This explains the frequency generation process when composing layers, highlighting the importance of input layer initialization for improving capacity. Remarkably, the amplitudes of these frequencies are given by complex functions of the hidden weights, introducing a challenging training. To address this, we prove that those are bounded by a term depending only on the hidden weights (Thrm 2). We apply this result to control the bandlimit of the INR during training. Some of TUNER’s results are showcased in Fig. 1, in comparison to previous work, illustrating substantial enhancements to sinusoidal INR modeling. In summary, our contributions are:
  • We introduce a novel trigonometric identity that expands any hidden neuron into a wide sum of sines with frequencies being integer combinations of the input frequencies (Fig 2). The corresponding amplitudes, determined by the hidden weights, have a useful upper bound, offering effective tools for controlling the resulting signal.
@@ 2 Related works
Implicit neural representations [30, 25, 8, 21, 32, 20] are a trending topic in machine learning, used to learn highly detailed signals in low-dimension domains. Current INR architectures use Fourier feature mappings [28] or sinusoidal activation functions [26] to bypass the spectral bias [17] common in ReLU multilayer perceptrons. The high representation capacity of sinusoidal INRs has motivated their use to represent a wide range of media objects. Examples include audio [6, 26], images [4, 27], face morphing [22], signed distance functions [23, 10, 26, 31, 13], displacement fields [33], surface animation [12, 14], multi-resolution signals [19, 10, 5, 29, 16], among others. Most of these exploit the derivatives of sinusoidal INRs in the loss functions.
@@ 3 Sinusoidal MLPs as Fourier series
Sinusoidal MLPs demonstrated high representational capacity with only a few hidden layers [26, 35, 19]. To understand how layer composition encodes this capacity, we propose to thoroughly investigate the structure of a $3$-layer sinusoidal MLP $f\!:\!\mathbb{R}^{d}\!\to\!\mathbb{R}$. For simplicity, we assume the codomain to be $\mathbb{R}$, however the same analysis holds for dimension $>\!\!1$. More precisely, we consider $f(\textbf{x})\!=\!\textbf{C}\!\circ\!\textbf{S}\!\circ\!\textbf{D}(\textbf{x})\!+\!e,$ with $\textbf{D}(\textbf{x})\!=\!\sin(\omega\textbf{x}\!+\!\varphi)$ being the input layer that projects x into a list of harmonics (input neurons) ${D}_{i}(\textbf{x})\!=\!\sin\left(\omega_{i}\textbf{x}+\varphi_{i}\right)$ with frequencies $\omega\!=\!(\omega_{1},\ldots,\omega_{m})\!\in\!\mathbb{R}^{m\times d}$ and shifts $\varphi\!\in\!\mathbb{R}^{m}$. Layer D is then composed with $\textbf{S}(\textbf{x})\!=\!\sin(\textbf{W}\textbf{x}+\textbf{b})$, where $\textbf{W}\in\mathbb{R}^{n\times m}$ is the hidden matrix, and $\textbf{b}\!\in\!\mathbb{R}^{n}$ the bias. Finally, $\textbf{C}\cdot\textbf{x}+e$ is an affine transformation with $\textbf{C}\in\mathbb{R}^{n\times 1}$ and $e\in\mathbb{R}$.
@@
Each hidden neuron $h_{i}$ of a 3-layer sinusoidal MLP has an amplitude-phase expansion of the form
@@
The Bessel functions $J_{k}$ appear in the Fourier series of $\sin\big{(}a\sin(x)\big{)}$ [2, Page 361] and Thrm 1 generalizes this result. See the proof and generalization in the Supp. Mat. Yüce et al. [34] present a similar expansion for neurons activated by polynomials, however, they do not derive an amplitude formula. Their expansion for sinusoidal neurons is restricted to a 3-layer MLP without bias, a critical parameter for Fourier representation. Ours incorporates bias and generalizes to deep MLPs.
@@ 4.1 Initialization as spectral sampling
Figure 4: Uniform init. of $\omega$ (top) and ours (bottom). Grayscale bands show INR’s gradient. Ours offers better signal/gradient reconstruction w/o gradient supervision. The MLPs $(m\!=\!128,n\!=\!256)$ with bandlimit $\mathscr{b}\!=\!87$ were trained for $3000$ epochs.
@@ 4.2 Bounding and hidden layer initialization
We use these bounds to initialize W such that each column is normal distributed with mean 00 and standard deviation $\tfrac{\mathscr{c}_{*}}{3}$, with $\mathscr{c}_{*}$ being the bound corresponding to that column. Note that SIREN [26] initializes $\textbf{W}\in\mathbb{R}^{n\times m}$ uniformly in $\small{\big{(}\!\!-\!{\sqrt{{6}/{m}}},{\sqrt{{6}/{m}}}\big{)}}$, and as $\small{\left|\textbf{W}\right|_{\infty}\!\leq\!\sqrt{{6}/{m}}<2}$ for $m>1$, it is bandlimited only at the beginning of training, since W is not controlled in any way during the optimization. Conversely, our initialization controls the bandlimit during all training and has faster convergence.
Additionally, we present a scheme to learn the bounding parameters during training. For this, we bound each $j$-column of the hidden matrix W using a $\tanh$ activation followed by multiplication with a trainable bounding parameter $\mathscr{c}_{j}$. Clearly, the resulting column is bounded by $\mathscr{c}_{j}$. In other words, we replace the hidden layer $\sin(\textbf{W}\textbf{x}+\textbf{b})$ by
@@ 5 Experiments
This section presents experiments regarding the initialization and bounding of frequencies of an INR $f$ considering the images from Kodak dataset [1] resized to $[-1,1]^{2}$. We describe the architecture of $f$ with hidden matrix $\textbf{W}\!\in\!\mathbb{R}^{n\times m}$ with the parameters $m,n$. We initialize $f$ following the scheme in Sec 4.1 and optimize it using Adam [9] with learning rate $10^{-4}$. For the quantitative evaluations, we report the mean and standard deviation of the Peak Signal-to-Noise Ratio (PSNR) evaluated on the test set ($10\%$ random pixels). To show that TUNER adds high order regularity to the reconstruction, we compare the analytical gradient $\nabla f$ to the Sobel filter of the ground truth. All experiments were run in a 12 GB NVIDIA GPU (TITAN X Pascal).
@@ 5.1 Initialization
First, we compare our initialization of $\omega$ ($36.2$dB) against the uniform distribution ($32.2$dB). We obtain a $4$dB improvement of PNSR on average (with a std of $1.6$dB). We also present an ablation over the parameter $\mathscr{l}$ which splits the square $[-\mathscr{b},\mathscr{b}]^{2}$ in regions of low (L) and high (H) frequencies. Here we use $\mathscr{b}=85$ and $\mathscr{l}\!=\!21$, $42$, $64$, and sample $\omega$ with $30\%$, $50\%$, and $70\%$ of coordinates in H. Table 1 shows the performance of the INR.
[…]
@@
Figure 8: Reconstructions with big $\mathscr{l}$ present overfitting on the $10\%$ unsupervised points. Gradient on the upper right corner.
@@ 5.2 Bounding
Fig 9 compares the reconstructions of an image of size $512^{2}$ wo/w our fixed bounding scheme. To demonstrate the impact of our scheme, we sample the input frequencies using $\mathscr{b}\!=\!256$ and $\mathscr{l}\!=\!100$. Without bounding (left), we effectively reconstruct the signal at the supervised pixels (30.8 dB), but with a noisy gradient (15.8 dB). That is, we are overfitting the signal. Conversely, bounding the hidden weights significantly improves both the signal (35.1 dB) and its gradient (27.8 dB). Our bounding scheme acts like a filter in both signal and gradient reconstruction.
Figure 9: Comparison of training wo/w fixed bounds. Note that the use of bounds preserves high order information of the signal. Left side of the figures show the network gradients.
We compare quantitatively our bounding scheme over the dataset, varying the parameters $\mathscr{b}$ and $\mathscr{c}_{\textbf{H}}$. Table 2 shows an improvement in the signal for all cases. Furthermore, the unsupervised gradient of the image is better learned when using fixed bounds, showing an average increase of 9.6dB against the standard training. Also, the last two lines of the table show the impact of the parameters $\mathscr{c}_{\textbf{L}}$, $\mathscr{c}_{\textbf{H}}$ over the signal and gradient reconstructions. We consider training with $\mathscr{c}_{\textbf{L}}>\mathscr{c}_{\textbf{H}}$ since we are prioritizing low input frequencies in the spectrum generation. Otherwise, this may generate high frequencies, possibly bypassing the Nyquist limit and resulting in a worse gradient reconstruction.
@@
Table 2: Training with bounding improves the signal and grad. reconstruction. The choice of the fixed bound $\mathscr{c}_{\textbf{H}}$ greatly impacts the grad. fidelity. The std of the difference between each image signal (grad) quality wo/w bounding is $1.4$dB (1.6dB) on average. Evaluations performed on the test set ($10\%$ of pixels).
@@
Figure 10: Comparison of signal and gradient (grayscale) when training with fixed/learned bounds. The learned bounds adjust the INR spectrum resulting in better reconstructions. Each blue point represents a pair $\left(|\omega_{j}|_{\infty},\mathscr{c}_{j}\right)$, where $c_{j}$ is the trained bound of $\omega_{j}$.
As expected, training lowers the bounds for high frequencies (H) while maintaining or increasing the bounds of low frequencies (L). Thus, we achieve better reconstruction in both the signal and gradient. In Fig 10 (right), we note that low frequencies (left of the green line) have significant bounds, showing their importance during training. These schemes allow us to initialize with higher frequencies than $\nicefrac{{B}}{{3}}$, accelerating convergence. Finally, we performed an ablation study on the regularization term $\mathcal{L}_{reg}$ to show its importance. Training w/wo regularization yielded $28.5/27.4$dB for the RGB and $27.1/24.0$dB for the gradient.
@@ 5.3 Comparisons
We first compare our method (TUNER) with SIREN, which requires choosing a threshold parameter $\mathscr{b}$, [^2] Denoted as $\omega_{0}$ in [26]. a critical task since higher $\mathscr{b}$ implies noisy reconstructions. TUNER also uses $\mathscr{b}$ to control the frequency generation, however, it provides more stable training with better reconstruction. To demonstrate this, we perform a comparison by varying $\mathscr{b}$ and the number of input ($m$) and hidden ($n$) neurons. First, consider the case where we have no information about the signal bandlimit besides the Nyquist limit $B$. Fig 11 presents a comparison between SIREN (top) and TUNER (bottom) with $\mathscr{b}=B$. Note that SIREN produces noisy reconstructions, while ours learns faster and does not suffer from noise. This is due to our bounding scheme, which limits the generation of high frequencies and provides stable training (Sec 4.2). In contrast, SIREN initializes its parameters such that the spectrum is initially bounded but does not maintain this guarantee during training.
Figure 11: Comparing SIREN (top) and TUNER (bottom) when we only know the Nyquist limit (i.e, $\mathscr{b}\!=\!B$). Since $\mathscr{b}$ is big, SIREN introduces high frequencies at the beginning of training, resulting in very noisy reconstructions. Conversely, ours starts training low frequencies, converging faster and being robust to $\mathscr{b}$.
@@
Figure 12: Comparison with SIREN when $\mathscr{b}=40$. TUNER improves training convergence in both cases ($m<n$, $m>n$) and stability when $m>n$. INR gradient shown in grayscale.
Next, we numerically evaluate these experiments, providing a quantitative comparison with SIREN and BACON. We compare different architectures ($m$, $n$) and spectrum bandlimits $\mathscr{b}$, evaluating the results in terms of signal and gradient PSNR over the supervised (Train) and unsupervised (Test) pixels. Table 3 presents these evaluations.
@@
As expected, SIREN performs well with $\mathscr{b}=\frac{B}{3}$. In this case, TUNER is perceptually similar to SIREN but improves gradient reconstruction and consistently outperforms BACON. When $\mathscr{b}=171$, TUNER offers superior quality in both signal and gradient, showing greater robustness to noise and superior performance even with $\mathscr{b}\geq\nicefrac{{B}}{{3}}$. Finally, with $\mathscr{b}=256$, our method surpasses SIREN in all cases. For BACON, our method achieves better signal reconstruction, with comparable gradients. However, the following experiment shows that BACON’s reconstruction has visible defects in this scenario. Note that we do not supervise the training with gradient information.